OpenStudy (anonymous):

Can someone please help me with a few "Solving Polynomial Equations" problems? Any help will be greatly appreciated! :D

OpenStudy (anonymous):

Here's the first one: Determine the zeros of f(x) = x3 – 3x2 – 16x + 48.

OpenStudy (anonymous):

@amistre64 @phi

OpenStudy (phi):

did you try plotting it in geogebra ?

OpenStudy (anonymous):

we have to show our work... :(

OpenStudy (phi):

What approach do they expect you to use to solve this ?

OpenStudy (anonymous):

The zeros of the function represent the values of x that make the function equal to zero. So you may replace the function notation with 0. f(x) = x2 + 7x + 6 0 = x2 + 7x + 6 Since the depressed equation is a quadratic equation, you may algebraically solve for x by factoring, using the quadratic formula, or by completing the square. Remember, looking at the discriminant is the best way to find out which method should be used. 0 = x2 + 7x + 6 a = 1, b = 7, c = 6 b2 – 4ac (7)2 – 4(1)(6) 49 – 4(1)(6) 49 – 24 25 The value of the discriminant is a perfect square, which means this quadratic may be solved by factoring. Review Show additional review on the discriminant f(x) = (x + 1)(x + 6) Set both factors equal to 0 and solve for x. 0=x + 1 0=x + 6 –1 –1 –6 –6 –1 = x –6 = x The two remaining zeros of f(x) = x3 + 3x2 – 22x – 24 are –1 and –6. These zeros may be tested using synthetic division. synthetic division with the factor, negative 1, in the top left corner, the coefficients 1, 3, negative 22, and negative 24 to the right, and 1, 2, negative 24, and 0 in the last rowsynthetic division with the factor, negative 6, in the top left corner, the coefficients 1, 3, negative 22, and negative 24 to the right, and 1, negative 3, negative 4, and 0 in the last row Both values of –1 and –6 have a remainder of 0. Therefore, they are both zeros of the function!

OpenStudy (anonymous):

That's what it says in the lesson

OpenStudy (phi):

yes, but you need at least one root (then you use synthetic division to reduce the problem to a quadratic) do they expect you to know the rational root test ?

OpenStudy (anonymous):

The rational root theorem?

OpenStudy (anonymous):

First, use the Rational Root Theorem to find the factors of p over q. The p value is the constant term –24. This is divisible by ±1, ±2, ±3, ±4, ±6, ±8, ±12, and ±24. The q value is the leading coefficient of 1. One is divisible only by itself and –1. Take each of the factors from p and write them as a fraction with the factors of q as the denominator. Make sure you account for all combinations. Factors of p: ±1, ±2, ±3, ±4, ±6, ±8, ±12, and ±24 Factors of q: ±1 Factors of p over q: ±1 over 1, ±2 over 1, ±3 over 1, ±4 over 1, ±6 over 1, ±8 over 1, ±12 over 1, ±24 over 1 These fractions may all be simplified so the factors of p over q are ±1, ±2, ±3, ±4, ±6, ±8, ±12, ±24. This means there are sixteen rational values that could possibly be zeros of the function. That's a lot of possibilities!

OpenStudy (anonymous):

This is the problem they're using f(x) = x3 + 3x2 – 22x – 24

OpenStudy (phi):

yes, that is it for your problem f(x) = x^3 – 3x^2 – 16x + 48 you need the factors of 48 divided by the factors of 1 (1 is the number in front of the x^3) the factors of 1 are just 1, and dividing by 1 does not change the answer. so your *possible factors* are the factors of 48 can you list the factors of 48 ?

OpenStudy (phi):

can you list the factors of 48 ?

OpenStudy (phi):

1 and 48 are the first set, so you put ±1, ±48 on the list can you find more factors ?

OpenStudy (phi):

does 2 divide into 48 ?

OpenStudy (anonymous):

+/-(1,2,3,4,6,8,12,24,48)

OpenStudy (phi):

yes, a very long list. now the idea is try each of those numbers in the equation f(x) = x^3 – 3x^2 – 16x + 48 replace x with a candidate number, do the arithmetic, and see if you get 0 once you find a number that gives 0, stop and we go on to the next step. there are 18 numbers (counting positive and negative), which is a boatload of work can you do it ?

OpenStudy (anonymous):

Since the degree of the polynomial is 3, doesn't that mean that there will be 3 zeros?

OpenStudy (phi):

yes, but once we find the first one, we can use synthetic division to make the problem a quadratic (degree 2), and then factor it. that will be less work

OpenStudy (anonymous):

okay. So I just plug in the factors until I find one that works?

OpenStudy (phi):

yes, it is not my idea of fun. start with ±1,±2±3 (assuming the answer will be a small number )

OpenStudy (anonymous):

3 works

OpenStudy (phi):

yes, so (x-3) is one of the factors and x=3 is one of the roots (makes the equation 0)

OpenStudy (anonymous):

okay, now what?

OpenStudy (phi):

do you know how to do synthetic division

OpenStudy (anonymous):

yeah

OpenStudy (phi):

the setup is